AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 398     Entry time: Fri May 12 16:24:20 2023
Author: RSS, TD 
Subject: V-I test 

Since yesterday evening, the bottom right detector was at -150 V. In the morning, the current was quite stable ~ 6.77 uA. The bias was switched off, the foils opened, both of the detector arms were moved in and out three times, and the chamber was tightly covered again. A negative bias voltage was applied from 0 to 150 V in steps of 10 V one by one to all four detectors, and the V-I results can be seen in Attachments 1-4. All detectors seem to be fine. The variation of the current with time at a constant bias voltage of -150 V was checked for all detectors and the results are in Attachments 5-8. Also, with an applied bias voltage of -150 V to the bottom right and left detector, the detector arms were moved in and out three times which showed a variation of 1-2 nA and 1 nA in the leakage current. For further conditioning of the bottom right detector, it is again left with a negative bias voltage of 150 V over the weekend.

Attachment 1: detector_1.png  18 kB  | Hide | Hide all
detector_1.png
Attachment 2: detector_2.png  19 kB  | Hide | Hide all
detector_2.png
Attachment 3: detector_3.png  19 kB  | Hide | Hide all
detector_3.png
Attachment 4: detector_4.png  18 kB  | Hide | Hide all
detector_4.png
Attachment 5: detector_1_variation.png  24 kB  | Hide | Hide all
detector_1_variation.png
Attachment 6: detector_2_variation.png  26 kB  | Hide | Hide all
detector_2_variation.png
Attachment 7: detector_3_variation.png  25 kB  | Hide | Hide all
detector_3_variation.png
Attachment 8: detector_4_variation.png  23 kB  | Hide | Hide all
detector_4_variation.png
ELOG V3.1.4-unknown